40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2820     Entry time: Tue Apr 20 18:02:22 2010
Author: Jenne 
Type: Update 
Category: COC 
Subject: New SRM and PRM Hung 

[Jenne, Steve]

We removed the old SRM and PRM from their cages, and are temporarily storing them in the rings which we use to hold the optics while baking.  Steve will work on a way to store them more permanently.

We then hung the new SRM (SRMU03) and new PRM (SRMU04) in the cages.  We were careful not to break the wires, so the heights will not have changed from the old heights.

The optics have not been balanced yet.  That will hopefully happen later this week.

ELOG V3.1.3-